June 21, 2013 Installation

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

Friday, June 21, 2013 at 1:30 PM At Jefferson Lab, the meeting will be held in F326


To connect from the outside, please use ESNET

  1. ) ESNET: 8542553
  2. )To connect by telephone (supported on request only) dial:
  • You can look up toll-free number at http://www.readytalk.com/intl
  • US and Canada: (866)740-1260 (toll free)
  • International: (303)248-0285 (toll call)
  • enter access code followed by the # sign: 3421244#

Desktop Sharing

You can view the computer desktop in the meeting room at JLab via the web.

  1. Go to http://esnet.readytalk.com
  2. In the "join a meeting" box enter the Hall D code: 3421244
  3. Fill in the participant registration form.

Up to GlueX/Hall-D Installation Meetings

Action Items From Last Meeting


  1. Announcements
  2. Updates to the Mechanical Installation Schedule ( Latest Rendition of Install Schedule) (Tim)
  3. Electronics Installation Schedule (Fernando) [| Schedule]
  4. Summer manpower
  5. Magnet Status
  6. Tagger Magnet Installation
  7. Microscope Installation
  8. Tagger Array Installation
  9. Pair Spectrometer Installation
  10. BCAL Installation
  11. FCAL Installation
  12. TOF Installation
  13. FDC Installation
  14. CDC Installation
  15. Slow Control Installation
  16. Trigger and DAQ Installation
  17. A.O.B.


Present: CMU (Curtis), JLab (Mark, Tim, Eugene, Elton, John, Hovannes, Yi, Elliott, Lubomir), UConn (James), IU (Matt)


Tim: New version of the mechanical installation schedule is available. Reflects changes due to extended schedule of magnet commissioning and repairing vacuum leaks.

Elliott: networking in the hall goes in as soon as possible. In the hall networking will be probably be installed in October.

Summer Students: Student labor has been very helpful in moving many tasks forward.

Magnet Status (Eugene): Refrigerator cleaning complete by next week and paperwork needs to be turned in. Once tests start then only a few weeks of testing are available to do tests with refrigerator. The current plan is to go to 1350 A for a brief period with a dewar attached to the refrigerator. The dewar would last only 12 hours. Then we go to 1300 A and map the field there.

The current plan is to refurbish the refrigerator, but this would not be enough cooling capacity for stable operation at 1350 A. The collaboration needs to decide what exactly is given. Current thinking is that the refurbished refrigerator is capable of sustaining stable operations at 1300 A.

Tagger Magnet (Tim): Vacuum chamber installed and leak tight and has been pumped down. Mapping figure has been completely designed. Planning to map in mid-late July and a student is doing the programming.

Microscope (James): no updates

Tagger Array: Tim: working with Catholic on installation details.

Pair Spectrometer: Tim: frame has been designed and needs to be constructed. Liping would like to assemble the low granularity portion of the tagger this summer.

BCAL (Elton): Ten modules remain to have light guide installation. Main focus is doing checks of dimensions of light guides. About 58 light guides need to be replaced in order to complete the installation. There is possible impact on schedule if USM does not deliver quickly. There are not enough LED boards presently to completely instrument 49 modules. Many LED boards were broken during light guide rework - those can be repaired in house. There is a question if spares should be consumed to instrument the 49th (spare) module or should be saved for service during operation. Student have been helping prepare the readout electronics.

FCAL (John,Elton): Darkroom frame is installed. Glass is stacked. Some environmental controls have arrived. Prototyping of the environmental system is underway. All monitoring boards were shipped from Athens and arrived at the end of the month.

TOF (Mark): Chuck has completed drawings for the FSU machine shop. Chuck has completed detailed drawing for the entire frame. Next week will have a dedicated installation and assembly during the TOF meeting.

FDC (Lubomir): Current tests are allowing opportunity to learn how to work with full cabling scheme and learn about interferences. Some trouble existed with cooling manifolds, but that has been fixed.

CDC (Curtis): Schedule remains unchanged. Exploring cooling.

Agenda additions: Slow controls, trigger and DAQ